OpenStudy (anonymous):

It takes Jake 9 hours longer than Joe to rebuild an engine. Working together, they can do the job in 20 hours. How long would it take each, working alone, to rebuild an engine?

OpenStudy (amistre64):

Jake = j+9 j + j+9 = 20

OpenStudy (amistre64):

2j = 11 j = 11/2 = 5.5 hours .... this doesnt look likeits gonna work out lke i thought lol

OpenStudy (anonymous):

let jack do in a hr. and joe in a+9 hr. in one hr both they do 1/a+1/(a+9). so 20(1/a +1/(a+9))=1 will give you a.

OpenStudy (amistre64):

ahhh.... i didnt do reciprocals lol

OpenStudy (anonymous):

darn cuz i had the same answer as amistre....so what the answer if that doesnt work out?